SlideServe
Menu
Préférences
© SlideServe · v2026

Probability Calculations with Angie Cook - Independent Events

harva avatar
8 diapositives · 227 Views · Sep 06, 2014

Learn about independent events and how to calculate marginal, joint, and conditional probabilities using Angie Cook's problem-solving approach. Solve Problem 2-16 with red, green, and white chips in an urn.

Probability Calculations with Angie Cook - Independent Events - slide 1
01 / 8Probability Calculations with Angie Cook - Independent Events
1 / 17

Télécharger la présentation

Probability Calculations with Angie Cook - Independent Events
Une image / un lien ci-dessous est fourni (tel quel) pour télécharger la présentation

Politique de téléchargement : le contenu du site vous est fourni TEL QUEL pour votre information et votre usage personnel ; il ne peut être vendu, concédé sous licence ni partagé sur d’autres sites sans l’accord de son auteur.Le contenu vous est fourni TEL QUEL, pour votre information et votre usage personnel uniquement. Téléchargez la présentation en cliquant sur ce lien. Lors du téléchargement, si vous ne parvenez pas à télécharger une présentation, l’éditeur a peut-être supprimé le fichier de son serveur.Pendant le téléchargement, si vous ne parvenez pas à obtenir une présentation, le fichier a peut-être été supprimé par l’éditeur.

F I N

Plus de contenu associé

Transcription de la présentation
  1. Independent Events That the occurrence of one event has no effect on the probability of occurrence of the second event. Angie Cook

  2. Probabilities: Independent Events • Marginal probability – the probability of an event occurring: P(A) • Joint probability – the probability of multiple, independent events, occurring at the same time: P(AB) = P(A)*P(B) • Conditional probability – • The probability of event B given that event A has occurred: P(B/A) = P(B) • Or, the probability of event A given that event B has occurred: P(A/B) = P(A) Angie Cook

  3. Problem 2-16 An urn contains 8 red chips, 10 green chips, and 2 white chips. A chip is drawn and replaced, and then a second chip drawn. What is the probability of …… Angie Cook

  4. P(R) 8/20 = .40 P(G) 10/20 = .50 P(W) 2/20 = .10 1.00 Angie Cook

  5. a) A white chip on the first draw? Marginal probability - P(A) P(W) = .10 Angie Cook

  6. b) A white chip on the first draw and a red on the second? Joint probability for two independent events – P(AB) = P(A)*P(B) P(WR) = P(W) * P(R) .10 * .40 = .04 Angie Cook

  7. c) two green chips being drawn? Joint probability for two independent events – P(AB) = P(A)*P(B) P(GG) = P(G) * P (G) .50 * .50 = .25 Angie Cook

  8. d) A red chip on the second, given that a white chip was drawn on the first? Conditional probability – for independent events P(AB) = P(B) P(R/W) = P(R) P(R/W) = .40 Angie Cook